# Quantum ergodicity and scrambling in quantum annealers

### v1

@@ Abstract
Quantum annealers play a major role in the ongoing development of quantum information processing and in the advent of quantum technologies. Their functioning is underpinned by the many-body adiabatic evolution connecting the ground state of a simple system to that of an interacting classical Hamiltonian which encodes the solution to an optimization problem. Here we explore more general properties of the dynamics of quantum annealers, going beyond the low-energy regime. We show that the unitary evolution operator describing the complete dynamics is typically highly quantum chaotic. As a result, the annealing dynamics naturally leads to volume-law entangled random-like states when the initial configuration is rotated away from the low-energy subspace. Furthermore, we observe that the Heisenberg dynamics of a quantum annealer leads to extensive operator spreading, a hallmark of quantum information scrambling. In all cases, we study how deviations from chaotic behavior can be identified when analyzing cyclic ramps, where the annealing schedule is returned to the initial configuration.
@@ I Introduction
Figure 1: (a) Schematic of a quantum annealing protocol. Two integrable Hamiltonians $H_{\rm M}$ and $H_{\rm P}$ are interpolated by chaotic Hamiltonian $H(s)$. Slowly varying the parameter $s$ from 00 to $1$ will connect the ground state of $H_{\rm M}$ to that the ground state of $H_{\rm P}$. (b) The forward ramp from $s(t=0)=0$ to $s(t=T)=1$ leads to the unitary map $U(T)\equiv U_{\rm FW}$. Ramping back from $s=1$ to $s=0$ generates the backwards unitary $U_{\rm BW}$ such that the cyclic process is given by $U_{\rm cycle}=U_{\rm BW}U_{\rm FW}$. (c) Half-chain entanglement entropy of the state $\left|\psi(T)\right>$ at the end of a forward ramp under annealing dynamics, for the Ising model described in Sec. II. The initial state is the spin coherent state $\left|\psi(0)\right>=\left|\theta=0,\varphi\right>$, where all qubits are pointing along the direction $(\cos(\varphi),\sin(\varphi),0)$ on the Bloch sphere. For $\varphi\rightarrow 0,\>\pi$, the dynamics connects the ground state of ±$\pm H_{\rm M}$ to that of ±$\pm H_{\rm P}$ and generate little entanglement. For intermediate values of $\varphi$, the final states acquire volume-law entanglement entropy, similar to that of random states.
@@
The emergence of chaotic properties in the dynamics of a quantum annealer can be readily observed in a simple example shown in Fig. 1c. There, we plot the half-chain entanglement entropy the state $\left|\psi(T)\right>$ at the end of the annealing schedule, as a function of the mean energy $\langle H_{\mathrm{M}}\rangle$ of the initial product state for different system sizes $N$. For $\langle H_{\mathrm{M}}\rangle=\pm N$, the initial state is the ground state of ±$\pm H_{\rm M}$, i.e. the standard annealing scenario, and in consequence we obtain that the corresponding final states are only slightly entangled. However, as the initial state is rotated away from the ground state (by angle $\varphi$, as described in Sect. III) we find that the final configuration acquires a entanglement entropy which grows with system size. At the peak, the final states exhibit clear volume-law entanglement entropy, well approximated by Page’s value [18]. We thus observe that the annealing dynamics is generating random-like states with very high entanglement, akin to what is expected from the time evolution of a fully quantum chaotic system.
@@
As we show in this paper, the observed chaotic features turn out to be generic and emerge in the dynamics of most initial states. Interestingly, we find that significant deviations from chaoticity are observed if one considers a cyclic annealing schedule, where the interpolation parameter $s(t)$ is driven from $s(t=0)=0$ to $s(t=T)=1$ and then back to $s(t=2T)=0$; see Fig. 1 (b). For such case, we observe the existence of eigenstates of $U(2T)$ that show clear signatures of low-energy states, in particular displaying area-law entanglement, and that the dynamics of global operators show unscrambling features.
@@
The rest of the manuscript is organized as follows. In Sec. II we introduce the model we use in this paper, characterize the onset of chaos and integrability for different parameter values, and discuss the connection between adiabaticity and the spectral gaps throughout the energy spectrum. In Sec. III we analyze the evolution of initial states that are rotated away from the ground state of $H_{\mathrm{M}}$. Expanding on the example discussed in Fig. 1 and considering both forward and cyclic adiabatic ramps. In Sec. IV we study properties of the many-body time-evolution operator $U(t)$ describing the dynamics. We focus on the analysis of the statistical properties of its eigenphases and entanglement structure of its eigenstates. We connect the behavior of these quantities to the emergence of chaotic signatures in the system for both forward and cyclic driving protocols. In Sec. V we study the adiabatic dynamics of the system in the Heisenberg picture. For that, we consider different sets of initial operators and probe the emergence of quantum information scrambling by tracking how the mean operator size evolves in time for the two driving protocols. Finally, in Sec. VI we summarize our findings, discuss how our results could inform new uses and applications of quantum annealers, and propose potential avenues for future work.
@@ II Physical model and driving protocol
Figure 2: Spectral properties of Hamiltonian $H(s)$ in Eq. (1). (a) Mean level spacing ratio $\overline{r}$, defined from Eq. (5), as a function of the interpolation parameter $s$. Average is taken over the bulk of the spectrum, here defined by not considering the top 5% lowest and highest energy levels. Dotted (dash-dotted) lines indicate the values of $\overline{r}$ for chaotic (GOE) and integrable (Poisson) models. (b) and (c) Spectral gaps as a function of $s$ for system sizes $N\in[8,12]$ (light to dark tones). Figure (b) shows the ground state gap $\Delta_{0}(s)$, while (c) shows the average gap $\sum_{j}\Delta_{j}/(\tilde{d}-1)$ with $\tilde{d}$ the relevant Hilbert space dimension. (d) Scaling of the minimum ground state gap and average gap with system size $N$; minimum is taken over $s\in[0,1]$. All calculations are performed in the positive parity sector of the Hamiltonian $H(s)$, whose dimension is $d=2^{N-1}$.
@@ II.1 Gap characterization and adiabaticity
We are interested in analyzing the dynamics of the system when the parameter $s=s(t)$ is slowly changed in time. We consider two protocols: a standard forward ramp, and a cyclic ramp. In the forward ramp, $s(t)=t/T$ and the Hamiltonian is interpolated between $H_{\rm M}$ at $t=0$ to $H_{\rm P}$ at $t=T$. This is the usual scenario in quantum annealing where one expects to connect the ground state of $H_{\rm M}$ to that of $H_{\rm P}$. In the cyclic ramp, on the other hand, $s(t)=t/T$ for $t\in[0,T]$ and $s(t)=(2T-t)/T$ for $t\in(T,2T]$, such that the Hamiltonian returns to $H_{M}$ at the final time $t=2T$. In this case, the system initially prepared in one of the eigenstates of $H_{\rm M}$ will return to the same eigenstate at $t=2T$ if the evolution is truly adiabatic, having acquired an overall geometric phase in the process [28].
@@
$\overline{\Delta}(s)=\frac{1}{\tilde{d}-1}\sum\limits_{j=1}^{\tilde{d}}\Delta_{j}(s),$ (7)
where $\tilde{d}$ indicates the dimension of the positive parity subspace, $d\simeq 2^{N-1}$. Curves are shown for various system sizes $N\in[8,12]$, revealing that the ground state gap remains finite as the system is scaled up, while the average gap shrinks exponentially. This is confirmed by the data in Fig. 2 (d) which shows the scaling of the minimum values of $\Delta_{0}(s)$ and $\overline{\Delta}(s)$ with system size. These results show that in order to strictly meet the adiabaticity condition Eq. (6) for every energy level, the evolution time would have to grow exponentially with system size, making it practically unfeasible.
@@ III Evolution of spin coherent states
[…]
@@ III.1 Forward ramps
First, we consider the evolution of every state $|\psi(0)\rangle\in\mathcal{S}$ with a forward ramp and compute $S_{A}$ of the final state $|\psi(T)\rangle$. When $|\psi(0)\rangle=|\pi/2,\pi\rangle$ with $\langle H_{\rm M}\rangle/N=1$, this is the standard quantum annealing scenario. Naturally, the ground state of $H_{\rm P}$ is reached at the end of the ramp, as is witnessed by the value $S_{A}=\log(2)$ in Fig. 1c (orange dashed line). The opposing limit, where $\varphi=0$ and $\langle H_{\rm M}\rangle/N=-1$, can be regarded as an anneal between $-H_{\rm M}$ and $-H_{\rm P}$, for which the final state is a product state and so $S_{A}=0$, as also seen in the Figure.
For initial states away from the two extremal cases, we find that they reach final configurations with $S_{A}$ rapidly converging to the Page value of random pure states (colored dashed lines in Fig. 1c). Interestingly, states with $S_{A}$ closer to the Page value also have mean-energy $\langle H_{\rm M}\rangle\approx 0$, corresponding with the middle of the energy spectrum, an observation which will be further discussed in Sec. IV. Finally, since $\langle\theta,\varphi|H_{\rm M}|\theta,\varphi\rangle=N\sin(\theta)\cos(\varphi)$, all other $|\theta,\varphi\rangle\notin\mathcal{S}$ have a mean-energy which rapidly decreases away from $\theta=\pi/2$. Hence the fate of the majority of states, $|\theta,\varphi\rangle$, in the manifold of SCSs, with the exception of small neighborhoods in the vicinity of $(\theta,\varphi)=(\pi/2,0),(\pi/2,\pi)$, i.e., $\langle H_{\rm M}\rangle/N=\pm 1$, is to be transported by the forward ramp into highly entangled states, displaying an entanglement entropy which grows with system size, and with some of them resembling random pure states.
@@ III.2 Cyclic ramps
Finally, we stress that the results discussed in Sec. III.1 and Sec. III.2 are not unique to product states; in fact, entangled initial states display a similar behavior. In App. A we report results for a set of initial states given by the eigenstates of $H_{\rm M}$, i.e., Dicke states, which are in general entangled. This suggests an initial-state-independet character of the reported behavior, connected to universal properties of the unitary operator $U(T_{\rm tot})$. We explore this aspect of the problem in the following section.
Figure 4: Spectral properties of the adiabatic unitary operator for both forward and cyclic ramps. (a) Mean level spacing ratio (MLSR) as a function of time for $0\leq t\leq 2T$ covering both forward and backward ramps. The green star indicates the first time instant $t^{\ast}$ at which the unitary becomes fully chaotic ($s\simeq 0.14$ for $N=14$), and the inset shows the behavior of this quantity with system size. The dash-dot line shows the MLSR for an integrable system and the dotted line that of a chaotic system. The bottom panels shows $s(t)$, with the forward ramp stopping at $t/T=1$ and the cyclic ramp defined over the full range of $t/T$. (b) Half-chain entanglement entropy of all the eigenstates of $U(T)$ as a function of their mean energy with respect to the problem Hamiltonian $H_{\rm P}$. (c) Half-chain entanglement entropy of all the eigenstates of $U(2T)$ as a function of their mean energy with respect to mixer Hamiltonian $H_{\rm M}$. Both in (b) and (c) the dots are colored based on the local density, with black indicating very low density and light orange indicating very high density, the dashed lines (bottom to top), indicate $\mathcal{S}=0,\log(2),0.5\log(N)-0.5$, respectively. In all panels the parameters are: $N=14$, $T=600$, $dt=0.5$.
@@ IV Chaos in the unitary evolution operator
To study the structure of the eigenstates, we use the half-partition entanglement entropy $S_{A}$ as defined un Eq. (12). For many-body systems in one dimension the half-chain entanglement entropy of individual eigenstates as a function of system size displays two dominant behaviors: i) Area-law entanglement with $S_{A}(N)\simeq K$ with $K$ a constant independent of the system size, and ii) volume-law entanglement with $S_{A}(N)=c_{1}N+c_{2}+O(N^{-1})$ a linear dependence on the system size. The coefficients $c_{1,2}$ allow us to distinguish between random gaussian states [39, 40] and random pure states which saturate the Page value [18], with $c_{1}=\log(2)/2$ and $c_{2}=-1/2$. Importantly, chaotic eigenstates of many-body Hamiltonians obey a volume-law of the half-chain entanglement entropy [41, 42, 13], thus one can diagnose the presence of this type of eigenstates based on the behavior of $S_{A}$.

### v2

@@ Abstract
Quantum annealers play a major role in the ongoing development of quantum information processing and in the advent of quantum technologies. Their functioning is underpinned by the many-body adiabatic evolution connecting the ground state of a simple system to that of an interacting classical Hamiltonian which encodes the solution to an optimization problem. Here we explore more general properties of the dynamics of quantum annealers, going beyond the low-energy regime. We show that the unitary evolution operator describing the complete dynamics is typically highly quantum chaotic. As a result, the annealing dynamics naturally leads to volume-law entangled random-like states when the initial configuration is rotated away from the low-energy subspace. Furthermore, we observe that the Heisenberg dynamics of a quantum annealer leads to extensive operator spreading, a hallmark of quantum information scrambling. In contrast, we find that when the annealing schedule is returned to the initial configuration (i.e. via a cyclic ramp), a subtle interplay between chaos and adiabaticity emerges, and the dynamics shows strong deviations from full ergodicity.
@@ I Introduction
Figure 1: (a) Schematic of a quantum annealing protocol. Two integrable Hamiltonians $H_{\rm M}$ and $H_{\rm P}$ are interpolated by chaotic Hamiltonian $H(s)$. Slowly varying the parameter $s$ from 0 to 11 will connect the ground state of $H_{\rm M}$ to that the ground state of $H_{\rm P}$. (b) The forward ramp from $s(t=0)=0$ to $s(t=T)=1$ leads to the unitary map $U(T)\equiv U_{\rm FW}$. Ramping back from $s=1$ to $s=0$ generates the backwards unitary $U_{\rm BW}$ such that the cyclic process is given by $U_{\rm cycle}=U_{\rm BW}U_{\rm FW}$. (c) Half-chain entanglement entropy of the state $\left|\psi(T)\right$> at the end of a forward ramp under annealing dynamics, for the Ising model described in Sec. II. The initial state is the spin coherent state $\left|\psi(0)\right>=\left|\theta=0,\varphi\right$>, where all qubits are pointing along the direction $(\cos(\varphi),\sin(\varphi),0)$ on the Bloch sphere. For $\varphi\rightarrow 0,\>\pi$, the dynamics connects the ground state of $\pm H_{\rm M}$ to that of $\pm H_{\rm P}$ and generate little entanglement. For intermediate values of $\varphi$, the final states acquire volume-law entanglement entropy, similar to that of random states.
@@
The emergence of chaotic properties in the dynamics of a quantum annealer can be readily observed in a simple example shown in Fig. 1c. There, we plot the half-chain entanglement entropy of the state $\left|\psi(T)\right$> at the end of the annealing schedule, as a function of the mean energy $\langle H_{\mathrm{M}}\rangle$ of the initial product state for different system sizes $N$. For $\langle H_{\mathrm{M}}\rangle=\pm N$, the initial state is the ground state of $\pm H_{\rm M}$, i.e. the standard annealing scenario, and in consequence we obtain that the corresponding final states are only slightly entangled. However, as the initial state is rotated away from the ground state (by angle $\varphi$, as described in Sect. III) we find that the final configuration acquires an entanglement entropy which grows with system size. At the peak, the final states exhibit clear volume-law entanglement entropy, well approximated by Page’s value [18]. We thus observe that the annealing dynamics is generating random-like states with very high entanglement, akin to what is expected from the time evolution of a fully quantum chaotic system.
@@
As we show in this paper, the observed chaotic features turn out to be generic and emerge in the dynamics of most initial states. Interestingly, we find that significant deviations from full ergodicity are observed if one considers a cyclic annealing schedule, where the interpolation parameter $s(t)$ is driven from $s(t=0)=0$ to $s(t=T)=1$ and then back to $s(t=2T)=0$; see Fig. 1 (b). For such case, we observe the existence of eigenstates of $U(2T)$ that show clear signatures of low-energy states, in particular displaying area-law entanglement, and that the dynamics of global operators show unscrambling features (i.e. operator growth followed by operator shrinking). Thus, in this regime we find that driving does not enhance ergodicity, but rather leads to a subtle interplay where adiabatic and chaotic features coexist.
@@
The rest of the manuscript is organized as follows. In Sec. II we introduce the model we use in this paper, characterize the onset of chaos and integrability for different parameter values, and discuss the connection between adiabaticity and the spectral gaps throughout the energy spectrum. In Sec. III we analyze the evolution of initial states that are rotated away from the ground state of $H_{\mathrm{M}}$, expanding on the example discussed in Fig. 1 and considering both forward and cyclic adiabatic ramps. In Sec. IV we study properties of the many-body time-evolution operator $U(t)$ describing the dynamics. We focus on the analysis of the statistical properties of its eigenphases and entanglement structure of its eigenstates. We connect the behavior of these quantities to the emergence of chaotic signatures in the system for both forward and cyclic driving protocols. In Sec. V we study the adiabatic dynamics of the system in the Heisenberg picture. For that, we consider different sets of initial operators and probe the emergence of quantum information scrambling by tracking how the mean operator size evolves in time for the two driving protocols. Finally, in Sec. VI we summarize our findings, discuss how our results could inform new uses and applications of quantum annealers, and propose potential avenues for future work.
@@ II Physical model and driving protocol
Figure 2: Spectral properties of Hamiltonian $H(s)$ in Eq. (1). (a) Mean level spacing ratio $\overline{r}$, defined from Eq. (5), as a function of the interpolation parameter $s$. Average is taken over the bulk of the spectrum, here defined by not considering the top 5% lowest and highest energy levels. Dotted (dash-dotted) lines indicate the values of $\overline{r}$ for chaotic (GOE) and integrable (Poisson) models. (b) and (c) Spectral gaps as a function of $s$ for system sizes $N\in[8,12]$ (light to dark tones). Figure (b) shows the ground state gap $\Delta_{0}(s)$, while (c) shows the average gap $\sum_{j}\Delta_{j}/(\overline{d}-1)$ with $\overline{d}$ the relevant Hilbert space dimension. (d) Scaling of the minimum ground state gap and average gap with system size $N$; minimum is taken over $s\in[0,1]$. All calculations are performed in the positive parity sector of the Hamiltonian $H(s)$, whose dimension is $d=2^{N-1}$.
@@ II.1 Gap characterization and adiabaticity
We are interested in analyzing the dynamics of the system when the parameter $s=s(t)$ is slowly changed in time. We consider two protocols: a standard forward ramp, and a cyclic ramp. In the forward ramp, $s(t)=t/T$ and the Hamiltonian is interpolated between $H_{\rm M}$ at $t=0$ to $H_{\rm P}$ at $t=T$. This is the usual scenario in quantum annealing where one expects to connect the ground state of $H_{\rm M}$ to that of $H_{\rm P}$. In the cyclic ramp, on the other hand, $s(t)=t/T$ for $t\in[0,T]$ and s(t)=(2Tt)/Ts(t)=(2T-t)/T for $t\in(T,2T]$, such that the Hamiltonian returns to $H_{M}$ at the final time $t=2T$. In this case, the system initially prepared in one of the eigenstates of $H_{\rm M}$ will return to the same eigenstate at $t=2T$ if the evolution is truly adiabatic, having acquired an overall geometric phase in the process [28].
@@
$\overline{\Delta}(s)=\frac{1}{\overline{d}-1}\sum\limits_{j=1}^{\overline{d}}\Delta_{j}(s),$ (7)
where $\overline{d}$ indicates the dimension of the positive parity subspace, $\overline{d}\simeq 2^{N-1}$. Curves are shown for various system sizes $N\in[8,12]$, revealing that the ground state gap remains finite as the system is scaled up, while the average gap shrinks exponentially. This is confirmed by the data in Fig. 2 (d) which shows the scaling of the minimum values of $\Delta_{0}(s)$ and $\overline{\Delta}(s)$ with system size. These results show that in order to strictly meet the adiabaticity condition Eq. (6) for every energy level, the evolution time would have to grow exponentially with system size, making it practically unfeasible.
@@ III Evolution of spin coherent states
$U(T_{\rm tot})\coloneq\mathcal{T}\left[e^{-i\int_{0}^{T_{\rm tot}}dtH(s(t))}\right]$ (8)
with $\mathcal{T}[.]$ the time ordering operator, $T_{\rm tot}=T,\>2T$ for the forward and cyclic ramps, respectively, and $s(t)$ chosen accordingly. In our numerical simulations we approximate the ideal protocol with a discretized version of the form
@@ III.1 Forward ramps
First, we consider the evolution of every state $|\psi(0)\rangle\in\mathcal{S}$ with a forward ramp and compute $S_{A}$ of the final state $|\psi(T)\rangle$. When $|\psi(0)\rangle=|\pi/2,\pi\rangle$ with $\langle H_{\rm M}\rangle/N=1$, this is the standard quantum annealing scenario. Naturally, the ground state of $H_{\rm P}$, given by (for $N$ even and open boundary conditions) $|{\rm GS}\rangle=\frac{1}{\sqrt{2}}(\lvert\uparrow\downarrow\rangle^{\otimes N/2}\pm\lvert\downarrow\uparrow\rangle^{\otimes N/2})$ with energy $E_{\rm GS}=-(N-1)$, is reached at the end of the ramp, as is witnessed by the value $S_{A}=\log(2)$ in Fig. 1c (orange dashed line). The opposing limit, where $\varphi=0$ and $\langle H_{\rm M}\rangle/N=-1$, can be regarded as an anneal between $H_{\rm M}$ and $H_{\rm P}$. Here the ground state of $H_{\rm P}$ given by $|{\rm GS}\rangle=\lvert\uparrow\rangle^{\otimes N}$, a product state with energy $E_{\rm GS}=-2N+1$  [^1] Notice that in absence of the longitudinal field $\sum_{i}\sigma_{i}^{z}$ the two states $|\uparrow\rangle^{\otimes N}$ and $|\downarrow\rangle^{\otimes N}$ are degenerate with energy E=(N1)E=-(N-1), thus the ground state of $H_{\rm P}$ corresponds to their even and odd superpositions. However, the inclusion of the longitudinal term shifts $|\uparrow\rangle^{\otimes N}$ down in energy to E=2N+1E=-2N+1 and $|\downarrow\rangle^{\otimes N}$ up in energy to $E=1$ effectively lifting the degeneracy, leaving $|\uparrow\rangle^{\otimes N}$ as the unique ground state., is reached at the end of the ramp, correspondingly $S_{A}=0$ as also seen in the Figure.
For initial states away from the two extremal cases, we find that the evolution reaches final configurations with $S_{A}$ rapidly converging to the Page value of random pure states (colored dashed lines in Fig. 1c). Interestingly, states with $S_{A}$ closer to the Page value also have mean-energy $\langle H_{\rm M}\rangle\approx 0$, corresponding with the middle of the energy spectrum, an observation which will be further discussed in Sec. IV. Finally, since $\langle\theta,\varphi|H_{\rm M}|\theta,\varphi\rangle=N\sin(\theta)\cos(\varphi)$, all other $|\theta,\varphi\rangle\notin\mathcal{S}$ have a mean-energy which rapidly decreases away from $\theta=\pi/2$. Hence the fate of the majority of states, $|\theta,\varphi\rangle$, in the manifold of SCSs, with the exception of small neighborhoods in the vicinity of $(\theta,\varphi)=(\pi/2,0),(\pi/2,\pi)$, i.e., $\langle H_{\rm M}\rangle/N=\pm 1$, is to be transported by the forward ramp into highly entangled states, displaying an entanglement entropy which grows with system size, and with some of them resembling random pure states.
@@ III.2 Cyclic ramps
Finally, we stress that the results discussed in Sec. III.1 and Sec. III.2 are not unique to product states; in fact, entangled initial states display a similar behavior. In App. A we report results for a set of initial states given by the eigenstates of $H_{\rm M}$, i.e., Dicke states, which are in general entangled. This suggests an initial-state-independent character of the reported behavior, connected to universal properties of the unitary operator $U(T_{\rm tot})$. We explore this aspect of the problem in the following section.
Figure 4: Spectral properties of the adiabatic unitary operator for both forward and cyclic ramps. (a) Mean level spacing ratio (MLSR) as a function of time for $0\leq t\leq 2T$ covering both forward and backward ramps. The green star indicates the first time instant $t^{\ast}$ at which the unitary becomes fully chaotic ($s\simeq 0.14$ for $N=14$), and the inset shows the behavior of this quantity with system size. The dash-dot line shows the MLSR for an integrable system and the dotted line that of a chaotic system. The bottom panels shows $s(t)$, with the forward ramp stopping at $t/T=1$ and the cyclic ramp defined over the full range of $t/T$. (b) Half-chain entanglement entropy of all the eigenstates of $U(T)$ as a function of their mean energy with respect to the problem Hamiltonian $H_{\rm P}$. (c) Half-chain entanglement entropy of all the eigenstates of $U(2T)$ as a function of their mean energy with respect to mixer Hamiltonian $H_{\rm M}$. Both in (b) and (c) the dots are colored based on the local density, with black indicating very low density and light orange indicating very high density, the dashed lines (bottom to top), indicate $\mathcal{S}=0,\log(2),0.5\log(N)-0.5$, respectively (see App. A for further information about the relevance of these values). all panels the parameters are: $N=14$, $T=600$, $dt=0.5$.
@@ IV Chaos in the unitary evolution operator
To study the structure of the eigenstates, we use the half-partition entanglement entropy $S_{A}$ as defined in Eq. (12). For many-body systems in one dimension the half-chain entanglement entropy of individual eigenstates as a function of system size displays two dominant behaviors: i) Area-law entanglement with $S_{A}(N)\simeq K$ with $K$ a constant independent of the system size, and ii) volume-law entanglement with $S_{A}(N)=c_{1}N+c_{2}+O(N^{-1})$ a linear dependence on the system size. The coefficients $c_{1,2}$ allow us to distinguish between random gaussian states [40, 41] and random pure states which saturate the Page value [18], with $c_{1}=\log(2)/2$ and $c_{2}=-1/2$. Importantly, chaotic eigenstates of many-body Hamiltonians obey a volume-law of the half-chain entanglement entropy [42, 43, 13], thus one can diagnose the presence of this type of eigenstates based on the behavior of $S_{A}$.